# spacing=nonfrench

### v2

@@ 4.3 RQ2: Latent Space Analysis
Although $|{\mathcal{D}_{\rm priv}}|$ is less important, $|{\mathcal{D}_{\rm infer}}|$ is capped by its size, and $|{\mathcal{D}_{\rm infer}}|$ has more impact on ${\mathcal{M}_{\rm L}}$’s accuracy. For example, on Path-MNIST with $\epsilon=10.0$, when $|{\mathcal{D}_{\rm priv}}|=1.5k$, LDPKiT-Rand’s accuracy range from 23.56% with $|{\mathcal{D}_{\rm infer}}|=15,500$ to 54.39% with $|{\mathcal{D}_{\rm infer}}|=500k$. Similarly, LDPKiT-Sup’s accuracy also improved from 42.99% to 76.79% when $|{\mathcal{D}_{\rm infer}}|$ increases. The most crucial requirement of $|{\mathcal{D}_{\rm infer}}|$ is the sufficient number of data points, e.g., ${\mathcal{M}_{\rm L}}$’s accuracy is always around 10% for SVHN when $|{\mathcal{D}_{\rm infer}}|$ is only 15,500, regardless of $|{\mathcal{D}_{\rm priv}}|$.
@@
For a more straightforward visualization of $|{\mathcal{D}_{\rm infer}}|$’s impact, we also present ${\mathcal{M}_{\rm L}}$’s accuracies with the setting of $|{\mathcal{D}_{\rm priv}}|=1.5k$ with varying $|{\mathcal{D}_{\rm infer}}|$ as bar graphs in Figure 8. Notably, a reasonably large ${\mathcal{D}_{\rm infer}}$ is required for LDPKiT-Sup to have an adequate accuracy (e.g., 250k for PathMNIST and 500k for Fashion-MNIST and SVHN). While increasing $|{\mathcal{D}_{\rm infer}}|$ enhances accuracy, the improvement becomes marginal once $|{\mathcal{D}_{\rm infer}}|$ exceeds 500k. For instance, ResNet-18 (${\mathcal{M}_{\rm L}}$) reaches 74.35%, 78.88% and 81.78% accuracies on ${\mathcal{D}_{\rm priv}}$ of Fashion-MNIST when it is trained on ${\mathcal{D}_{\rm infer}}$ with $\epsilon$ set to 2.0 and sizes of 500k, 1M and 2.2M, respectively. In other words, it is unnecessary to query all possible data points in ${\mathcal{D}_{\rm cand}}$, as discussed in Section 4.2. The trained ${\mathcal{M}_{\rm L}}$ with LDPKiT-Sup can achieve satisfactory accuracies using only about 20% of ${\mathcal{D}_{\rm infer}}$ for querying. Appendix D.3 presents the evaluation results for other $\epsilon$ values on the same benchmarks. The results match the conclusions drawn from Figure 8.
@@
It is important to note that the subsets of ${\mathcal{D}_{\rm infer}}$ used to train ${\mathcal{M}_{\rm L}}$ are randomly selected in this section of experiments. More advanced methods, such as active learning and core-set strategies [33, 26], could be employed to identify a smaller yet representative subset of ${\mathcal{D}_{\rm infer}}$, effectively approximating the full dataset’s ability to train the model or capture the dataset’s overall distribution. Exploring such data selection methods for ${\mathcal{D}_{\rm infer}}$ size reduction is left for future work.
4.5 Effect of Noise on Overfitting to ${\mathcal{D}_{\rm priv}}$ and Generalization to ${\mathcal{D}_{\rm val}}$ Across $\epsilon$ Values
LDP bounds privacy leakage rather than eliminating it. As defined in Definition 2.1, larger $\epsilon$ values reduce the LDP noise, weakening privacy protections. In reality, the user is interested in labelling ${\mathcal{D}_{\rm priv}}$. However, to safeguard privacy, she may choose to query only a subset of ${\mathcal{D}_{\rm priv}}$ while maintaining a reasonable $|{\mathcal{D}_{\rm infer}}|$ and label the rest of the data she never queried with the extracted ${\mathcal{M}_{\rm L}}$ model if it can generalize to unseen data points ${\mathcal{D}_{\rm val}}$ (i.e., also attributes to $|{\mathcal{D}_{\rm priv}}|$ insensitivity as discussed in Section D.3). To explore this, we analyze the impact of $\epsilon$ on generalization. Improved generalization implies better privacy, as the user can reduce $|{\mathcal{D}_{\rm priv}}|$ for generating queries while still extracting a model that generalizes to the remaining private data. Figures 9 and 10 show that ResNet-18 (${\mathcal{M}_{\rm L}}$) achieves higher accuracy on ${\mathcal{D}_{\rm priv}}$ than on ${\mathcal{D}_{\rm val}}$ when $\epsilon$ is high (weaker privacy protection), particularly with LDPKiT-Rand. As $\epsilon$ decreases, ${\mathcal{D}_{\rm val}}$ accuracy improves, eventually matching or exceeding ${\mathcal{D}_{\rm priv}}$ accuracy, indicating reduced overfitting and better generalization. Similar trends are observed with MobileNetV2 in Figures 14 and 15 in Appendix D.2. Therefore, the user may obtain better privacy protection by reducing $|{\mathcal{D}_{\rm priv}}|$ while she increases $\epsilon$. Further analysis can be performed to provide deeper insights. For instance, we can measure privacy leakage using metrics such as membership inference attack success rates and attribute inference metrics to help examine the relationship between $\epsilon$, overfitting and privacy leakage. We can also analyze the mutual information between ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm infer}}$ under different $\epsilon$ settings. We leave the analyses as future work.
5 Related work
Knowledge transfer techniques.
@@ 5 Related work
Knowledge distillation is a knowledge transfer technique that distills a large teacher model into a smaller student model while preserving model performance [34, 35, 36, 37, 38, 39, 40]. The conventional use case is model compression, enabling the model deployment in a resource-restricted environment. Knowledge can be transferred in different forms, such as logits, model parameters, intermediate layers’ activations or features, and their interrelationships [35]. Knowledge transfer also has adversarial applications. A model extraction attack is an adversarial example in which the attacker reproduces a model stealthily by stealing its parameters, decision boundaries, or functionalities. It demonstrates that an iterative query-based knowledge transfer process from a high-performance model can be performed via a prediction query interface [41, 42, 43]. Model extraction can be challenging without knowledge of the victim model’s training data distribution [44], successful model extractions with partial or zero knowledge of the victim model and training data [45, 42, 44, 46, 47] can be costly in queries and require more information than hard labels. Related defences [48, 49, 50, 51] and analyses [52, 53] are also in active research. Rather than model compression or extraction, we incorporate knowledge transfer for privacy preservation in a non-adversarial manner to recover the utility loss brought by LDP. Therefore, regarding the accuracies, our mechanism still demonstrates the privacy-utility trade-off rather than creating a competitive surrogate model that violates cloud services’ terms of use. We qualitatively and quantitatively demonstrate that LDPKiT differs from the model extraction attack in Section 6.1.
Noise injection and differential privacy (DP).
@@
DP can be used either locally [54] or globally [55, 56, 57], and both methods provide provable privacy guarantees. Global differential privacy (GDP) shares original input data with a trusted data curator, which then applies noise to the aggregated data. In this case, the curator has access to the original sensitive data. To remove this point of trust, LDPKiT uses a local differential privacy (LDP) mechanism. In LDP, the data source (i.e., the user) adds noise to each individual query before data transmission to the cloud, so she has full control over privacy protection. The side effect is that the noise is aggregated on the cloud, so LDP methods usually provide lower utility than GDP at the same level of privacy protection (i.e., noise level). While we add noise to original inputs before offloading inference to the cloud, similar to [15], noise can also be injected into inference frameworks deployed on a split computation setting [58, 59, 60, 61], where the DNN is partitioned between the cloud and edge devices. These schemes involve a white-box model, so noise can be added to intermediate representations, which is different from our setting. A common challenge of DP schemes is to find a balance between utility and privacy.
Other privacy protection techniques.
@@
One class of privacy protection methods is data encryption with homomorphic algorithms, which suffers from high computational overheads [12, 13]. In contrast, LDPKiT is more efficient, as it does not require complicated computation to be performed for each query. Hardware-assisted inference in Trusted Execution Environments (TEEs) is another approach [14]. A TEE is a secure area within a processor that provides a safe environment for sensitive code execution, preventing unauthorized access. Slalom puts the computation in a TEE to address inference privacy on remote services [14]. However, Slalom does not protect against the risks of side-channel attacks. Since TEEs have access to the original data, privacy breaches can still happen if the attackers compromise the TEEs [62, 63, 64]. Side-channels are not a threat for LDPKiT since it does not transmit the original data, and any privacy leakage is bounded by the LDP noise.
6 Discussion and Limitations
Table 4: Normalized Zest distance results with Cosine distance metric on ${\mathcal{M}_{\rm R}}$ and ${\mathcal{M}_{\rm L}}$.
$\epsilon$=2.0 $\epsilon$=1.5 $\epsilon$=1.5 $\epsilon$=1.25 $\epsilon$=10.0 $\epsilon$=7.0
ResNet-152 ResNet-18 LDPKiT-Sup 2.5598 3.2729 1.1948 1.1859 5.3133 5.4322
LDPKiT-Rand 3.6643 4.2645 1.6503 2.4100 6.1018 4.2305
MobileNetV2 LDPKiT-Sup 5.5795 5.1662 1.3259 1.2905 7.2874 7.2874
LDPKiT-Rand 3.1595 2.7800 1.6183 1.7758 7.2874 7.2874
6.1 Ethical Use of Model Extraction Principles
@@
While our approach involves knowledge transfer from a remote model to a local model, it differs from adversarial model extraction attacks [45, 42, 44, 46, 47]. The deployment of LDPKiT does not aim to replicate the economic value of commercial models or compete with model owners. Instead, LDPKiT ethically integrates model extraction principles to protect user inference data information, prioritizing data privacy over adversarial exploitation.
@@
Existing defenses against model extraction attacks often lack generalizability [19, 21], and advanced techniques [20] can bypass them, risking economic losses for commercial model owners. However, this is not a concern in our case. LDPKiT’s extraction objective focuses solely on protecting data privacy, not economic gain. At fairly privacy-protective noise levels, such as when SIDP drops to approximately 25%, the ${\mathcal{M}_{\rm L}}$ model generated by LDPKiT exhibits an accuracy trade-off for privacy ranging from 10% to 30%, depending on the noise level, model and dataset. Notably, ${\mathcal{M}_{\rm L}}$ is not competitive with ${\mathcal{M}_{\rm R}}$; for instance, ResNet-18 (${\mathcal{M}_{\rm L}}$) show accuracy drops of 12.96% and 21.23% on Fashion-MNIST with LDPKiT-Sup at $\epsilon$ values of 2.0 and 1.5, respectively. Given this lack of competitiveness, ${\mathcal{M}_{\rm L}}$ does not serve the objectives of adversarial model extraction and adheres to the non-competition terms of use outlined by major commercial model providers [18].
Furthermore, we assess the similarity between ${\mathcal{M}_{\rm L}}$ created by LDPKiT and ${\mathcal{M}_{\rm R}}$ in Section 4.2 using Zest distances [65]. Zest is a detection scheme for model extraction attacks that computes distances between two models based on LIME’s model-agnostic explanations [66]. We use Zest because of its architecture independence, the model’s black-box access requirement, and its perfect accuracy in model extraction detection with Cosine distance metric. Zest supports $l_{1}$, $l_{2}$, $l_{\infty}$ norm and Cosine distances. According to the authors of Zest, the method achieves 100% accuracy in detecting model extraction attacks when employing the Cosine distance metric [65], which is the metric presented in Table 4. Details of the detection procedure are outlined in Appendix E.1. We present the normalized Zest distances, $\overline{D_{z}}$, with the Cosine distance metric in Table 4. Results for other supported metrics are provided in Appendix E.2. As explained in Appendix E.1, an adversarial model extraction attack occurs when $\overline{D_{z}}<1$. Table 4 shows that LDPKiT does not contribute to model theft at any noise level since all $\overline{D_{z}}>1$. In conclusion, the Zest results confirm that the extracted ${\mathcal{M}_{\rm L}}$ model does not closely mimic ${\mathcal{M}_{\rm R}}$ and primarily serves to recover utility for ${\mathcal{D}_{\rm priv}}$. Combined with Section 4 results, these findings validate the ethical application of model extraction principles in LDPKiT, while upholding privacy objectives without compromising the benefits, rights or interests of any stakeholders.
6.2 Limitations and Future Work

### v3

@@ IV-C RQ2: Latent Space Analysis
$|{\mathcal{D}_{\rm infer}}|$ determines the cost and has more impact on $\mathcal{M}_{\rm L}$}’s accuracy. Although $\mathcal{D}_{\rm priv}$} is less critical in size, it ultimately limits the maximum size of $\mathcal{D}_{\rm infer}$} we can generate. For example, on PathMNIST with $\epsilon=10.0$, when $|{\mathcal{D}_{\rm priv}}|=1.5k$, LDPKiT-Rand’s accuracy range from 23.56% with $|{\mathcal{D}_{\rm infer}}|=15,500$ to 54.39% with $|{\mathcal{D}_{\rm infer}}|=500k$. Similarly, LDPKiT-Sup’s accuracy also improved from 42.99% to 76.79% when $|{\mathcal{D}_{\rm infer}}|$ increases. In contrast, $\mathcal{M}_{\rm L}$}’s accuracy is always around 10% for SVHN when $|{\mathcal{D}_{\rm infer}}|$ is only 15,500, regardless of $|{\mathcal{D}_{\rm priv}}|$. Therefore, having a sufficient number of data points in $\mathcal{D}_{\rm infer}$} is crucial.
@@
For a more straightforward visualization of $|{\mathcal{D}_{\rm infer}}|$’s impact, we also present $\mathcal{M}_{\rm L}$}’s accuracies with a fixed setting of $|{\mathcal{D}_{\rm priv}}|=1.5k$ and varying $|{\mathcal{D}_{\rm infer}}|$ as bar graphs in Figure 7. Notably, a reasonably large $\mathcal{D}_{\rm infer}$} is required for LDPKiT-Sup to have an adequate accuracy (e.g., 250k for PathMNIST and 500k for Fashion-MNIST and SVHN). While increasing $|{\mathcal{D}_{\rm infer}}|$ (i.e., higher querying cost) enhances accuracy, the improvement becomes marginal once $|{\mathcal{D}_{\rm infer}}|$ exceeds 500k. For instance, ResNet-18 ($\mathcal{M}_{\rm L}$}) achieves accuracies of 74.35%, 78.88% and 81.78% on $\mathcal{D}_{\rm priv}$} of Fashion-MNIST when trained on $\mathcal{D}_{\rm infer}$} with $\epsilon$ set to 2.0 and $|{\mathcal{D}_{\rm infer}}|$ set to 500k, 1M and 2.2M, respectively. In other words, it is unnecessary to query all possible data points in $\mathcal{D}_{\rm infer}$}, as discussed in Section IV-B. The trained $\mathcal{M}_{\rm L}$} with LDPKiT-Sup can achieve satisfactory accuracies using only about 20% of $\mathcal{D}_{\rm infer}$} for querying. Appendix A-D2 presents the evaluation results for other $\epsilon$ values on the same benchmarks, which match the conclusions drawn from Figure 7.
@@
It is important to note that the subsets of $\mathcal{D}_{\rm infer}$} used to train $\mathcal{M}_{\rm L}$} are randomly selected in this section of the experiments. Advanced methods, such as active learning and core-set strategies [42, 35], could be employed to identify a smaller yet representative subset of $\mathcal{D}_{\rm infer}$}, effectively capturing the dataset’s overall distribution to train the model. Exploring such data selection methods for $\mathcal{D}_{\rm infer}$} size reduction is left for future work.
Currently, there are no universal guidelines for estimating a precise lower bound on $\lvert{\mathcal{D}_{\rm priv}}\rvert$ that guarantees reliable performance. The effective lower bound depends on several factors, including (i) the complexity of the remote model and task (e.g., number of classes, class imbalance), (ii) the degree of noise introduced by the LDP mechanism, and (iii) the diversity and representativeness of the available private samples. In practice, an end user with limited data can gauge feasibility through small-scale pilot runs. For example, if $\lvert{\mathcal{D}_{\rm priv}}\rvert$ is sufficiently large, the user may expand $\mathcal{D}_{\rm infer}$} by 5%, 10%, or 20% and monitor the resulting accuracy trend. Developing systematic methods to estimate realistic thresholds for $\lvert{\mathcal{D}_{\rm priv}}\rvert$ remains an important direction for future research.
TABLE IV: Classification accuracy and image similarity under various $\epsilon$ settings. Similarity is computed between original and noisy images (before reconstruction), and between original and denoised images (after reconstruction). For reconstruction quality: $\uparrow$ indicates higher is better; $\downarrow$ indicates lower is better.
Dataset $\bm{\epsilon}$ Recon.
ResNet-152
Acc. (%)
$\uparrow$
MSE $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SVHN 2.0 Before 16.80 0.0681 0.1210 0.7591
After 14.20 0.1524 0.0825 0.7183
1.5 Before 15.30 0.0929 0.0850 0.7747
After 14.40 0.1436 0.0829 0.7301
1.25 Before 12.50 0.1098 0.0681 0.7826
After 13.00 0.1382 0.0810 0.7347
Fashion -MNIST 2.0 Before 16.10 0.0942 0.3869 0.5568
After 16.70 0.1257 0.3180 0.4492
1.5 Before 14.80 0.1123 0.3406 0.5928
After 16.30 0.1395 0.2953 0.5044
Path MNIST 10.0 Before 22.78 0.0132 0.4823 0.2237
After 23.89 0.0853 0.3870 0.2593
7.0 Before 17.56 0.0155 0.3936 0.2507
After 21.89 0.1012 0.3110 0.2805
IV-E Evaluation of Data Reconstruction Risks
To rigorously evaluate whether the amount of LDP noise added to each private image is sufficient, we audit the privacy guarantees of LDPKiT following the methodology of [31]. Specifically, we conduct data reconstruction attacks using diffusion models (i.e., DDPM [43]) that learned the image priors (i.e., trained on the original images). Our experiment spans three datasets: SVHN, Fashion-MNIST, and PathMNIST. For each dataset, we randomly select a subset of 1,000 images (900 for PathMNIST) processed with LDPKiT-Sup, and quantitatively measure the data reconstruction success with established theoretical bounds: mean squared error (MSE), VGG-based learned perceptual image patch similarity (LPIPS) [44, 45], and structural similarity index measure (SSIM) [46]. We report results across different $\epsilon$ values in Table IV. Our measurements show that the reconstructed images are generally more dissimilar from the original images after reconstruction (i.e., lower SSIM, higher LPIPS, and higher MSE), indicating that the reconstruction fails. We also show that a fully trained classifier (ResNet-152) is unable to classify the reconstructed images correctly. We show some sample reconstructed images in Figure 12 in Appendix A-D3.
V Related work
V-A Knowledge transfer techniques
@@
Knowledge distillation is a knowledge transfer technique that distills a large teacher model into a smaller student model while preserving model performance [47, 48, 49, 50, 51, 52]. The conventional use case is model compression, enabling the model deployment in a resource-restricted environment. Knowledge can be transferred in different forms, such as logits, model parameters, intermediate layers’ activations or features, and their interrelationships [48]. Knowledge transfer also has adversarial applications. A model extraction attack is an adversarial example in which the attacker reproduces a model stealthily by stealing its parameters, decision boundaries, or functionalities. It demonstrates that an iterative query-based knowledge transfer process from a high-performance model can be performed via a prediction query interface [53, 54, 55]. Model extraction can be challenging without knowledge of the victim model’s training data distribution [56]. Successful model extractions with partial or zero knowledge of the victim model and training data [57, 54, 56, 58, 59] may require more information than hard labels. Related defenses [60, 61, 62, 63] and analyses [64, 65] are in active research. Rather than model compression or adversarial extraction for the purpose of theft, we incorporate extraction techniques for privacy preservation in a non-adversarial manner to recover the utility loss brought by LDP. Our mechanism still demonstrates the privacy-utility trade-off rather than creating a competitive surrogate model that violates ML services’ terms of use.
V-B Noise injection and differential privacy (DP)
@@
DP can be used locally [66] or globally [67, 68, 69], and both provide provable privacy guarantees. Global differential privacy (GDP) shares original input data with a trusted data curator, which then applies noise to the aggregated data. In this case, the curator has access to the original sensitive data. To remove this point of trust, LDPKiT applies local differential privacy (LDP). In LDP, the data source (i.e., the user) adds noise to each individual query before data transmission to the ML service, so she has full control over privacy protection. The side effect is that the noise is aggregated on a server, so LDP methods usually provide lower utility than GDP at the same level of privacy protection (i.e., noise level) [70]. While we add noise to original inputs before offloading inference to the cloud, similar to [71], noise can also be injected into inference frameworks deployed on a split computation setting [72, 73, 74, 75], where the DNN is partitioned between the cloud and edge devices. These schemes involve a white-box model, so noise can be added to intermediate representations, which differs from our setting, where we assume the model provider is reluctant to share the model with the user.
V-C Other privacy protection techniquess
@@
One class of privacy protection methods is data encryption with homomorphic algorithms, which suffers from high computational overheads [15, 16]. For experiments on SVHN, LDPKiT begins to demonstrate lower computational overhead when the number of queries is around 800. Specifically, full knowledge transfer with LDPKiT on 800 data points only takes 20.73 seconds, compared to Lancelot [76], which takes 22.984 seconds for one inference on the same dataset, and OpenFHE [77], which requires significantly more time at 484.184 seconds. Furthermore, LDPKiT performs knowledge transfer once, enabling subsequent inference from the same sensitive data distribution on the trained local model with negligible overhead. In contrast, encryption schemes continue to impose per-query costs on the remote model, resulting in higher cumulative costs when the user wants to label a reasonably large dataset. Hardware-assisted inference in Trusted Execution Environments (TEEs) is an alternative approach [17]. A TEE is a secure area within a processor that provides a safe environment for sensitive code execution, preventing unauthorized access. Slalom puts the computation in a TEE to address inference privacy on remote services [17]. However, Slalom does not protect against the risks of side-channel attacks. Since TEEs have access to the original data, privacy breaches can still happen if the attackers compromise the TEEs [78, 79, 80]. Side-channels are not a threat to LDPKiT, since it does not transmit the original data, and any privacy leakage is bounded by the LDP noise.
V-D Federated Learning (FL)
FL has emerged as a popular approach to address privacy concerns by enabling collaborative model training without directly sharing users’ raw data with a central server [81, 82, 83, 84, 85]. However, FL primarily protects users’ training data, which is an orthogonal concern to our setting, and it does not mitigate privacy risks during inference. Moreover, FL is vulnerable to gradient inversion attacks that reconstruct local training data from gradient updates, which necessitates extra privacy protection [86, 87, 88, 89]. Finally, FL assumes that users possess labelled local datasets, which is not applicable in our case, where the objective is to annotate unlabelled data. Consequently, concerns regarding potential data reconstruction attacks and the integrity of the server imply that relying solely on using shared models from service providers like FL is insufficiently secure and privacy-preserving; therefore, FL is not an ideal solution in our scenario.
VI Discussion, Limitations and Future Work
VI-A Ethical Use of Model Extraction Principles
@@
While our approach involves knowledge transfer from a remote model to a local model, it differs from adversarial model extraction attacks [57, 54, 56, 58, 59]. LDPKiT ethically integrates model extraction principles to protect user inference data information, prioritizing data privacy over adversarial exploitation.
@@
Existing defenses against model extraction attacks often lack generalizability [25, 27], and advanced techniques [26] can bypass them, risking economic losses for commercial model owners. However, this is not a concern in our case. LDPKiT’s extraction objective focuses solely on protecting data privacy, not economic gain. At fairly privacy-protective noise levels, such as when SIDP drops to approximately 10% to 25%, the $\mathcal{M}_{\rm L}$} model generated by LDPKiT exhibits an accuracy trade-off for privacy ranging from 10% to 30%, depending on the noise level, model, and dataset. Notably, $\mathcal{M}_{\rm L}$} is not competitive with $\mathcal{M}_{\rm R}$}; for instance, ResNet-18 ($\mathcal{M}_{\rm L}$}) shows an accuracy drop of 12.96% on Fashion-MNIST with LDPKiT-Sup at $\epsilon=2.0$. Given this lack of competitiveness, $\mathcal{M}_{\rm L}$} does not serve the objectives of adversarial model extraction and adheres to the non-competition terms of use outlined by major commercial model providers [24].
In this work, our definitions of non-competitiveness, and non-stealthiness for ethical and non-adversarial extraction are scoped to the research setting: (i) Non-competitiveness means that the extracted $\mathcal{M}_{\rm L}$} models do not approach the utility or economic value of their $\mathcal{M}_{\rm R}$} counterparts and therefore cannot substitute or rival them in any commercial sense. (ii) Non-stealthiness indicates that LDPKiT does not attempt to evade detection or disguise its operation; its goal is explicitly protective, rather than adversarial. These working definitions emphasize that LDPKiT is designed and evaluated within a research context, to advance user data protection, rather than challenging or undermining existing commercial ecosystems. We further acknowledge that the notions of non-competitiveness and non-stealthiness require careful interpretation and, in practice, would necessitate formal reviews by legal and compliance experts.
VI-B Limitations and Future Work